# Center vortices and localized Dirac modes in the deconfined phase of 2+1 dimensional lattice $\mathbb{Z}_{2}$ gauge theory

### v1

@@ II.5 Sea/islands picture and Dirac-Anderson Hamiltonian
For the staggered Dirac operator the sea/islands picture briefly sketched in the Introduction can be understood more precisely in the language of the Dirac-Anderson Hamiltonian [31, 32, 24, 20]. In $d+1$ dimensions, in the basis where the temporal part of the staggered operator is diagonal, the “Hamiltonian” $-iD^{\mathrm{stag}}$ reads $H^{\rm DA}=-i\Omega^{\dagger}D^{\mathrm{stag}}\Omega$, where $\Omega$ is a suitable unitary change of basis and [20]
@@
${\cal A}_{\psi}(\lambda)\equiv\frac{\left\langle\sum_{l}\delta(\lambda-\lambda_{l})\sum_{\vec{x}}{\cal A}(\vec{x})\sum_{t=0}^{N_{t}-1}|\psi_{l}(t,\vec{x})|^{2}\right\rangle\,,}{\left\langle\sum_{l}\delta(\lambda-\lambda_{l})\right\rangle\langle{\cal A}(\vec{x})\rangle}\,.$ (30)
@@ III.1 Mobility edge
In Figs. 1 and 2 we show the normalized spectral density $\rho(\lambda)$, Eq. (10), and the average mode size $\overline{\mathrm{IPR}^{-1}}$, see under Eq. (20), along the low-lying spectrum. The spectral density is small but nonzero near $\lambda=0$, and starts to increase rapidly in the bulk, above $\lambda_{(0)}$. In the same region the mode size is essentially volume-independent, indicating localization of the eigenmodes. Peak structures are present in $\rho$ at specific values of $\lambda$ below $\lambda_{(0)}$, where one correspondingly finds dips in the mode size. We will return on this feature below in Sec. III.5.
We then studied the statistical properties of the unfolded spectrum. Since degenerate eigenvalues are absent in the spectral region and for the volumes under consideration, we have obtained the unfolded eigenvalues numerically by sorting in ascending order all the eigenvalues on all the configurations for a given choice of $\beta$ and $N_{s}$, and assigning to each eigenvalue its rank divided by the number of configurations. We then used them to numerically estimate $\overline{I}_{s_{0}}(\lambda,N_{s})$, Eq. (25), as explained above, and studied its volume dependence. The behavior expected for a BKT-type Anderson transitions is clearly visible in Fig. 3, where we show $\overline{I}_{s_{0}}$ for different spatial volumes at $\beta=0.734$ and $\beta=0.7425$ (here $\Delta\lambda=0.0125$). While the low modes tend towards Poissonian statistics as the system size increases, higher up in the spectrum the spectral statistics does not change with $N_{s}$. The lowest Matsubara frequency, $\lambda_{(0)}$, and the mobility edge, $\lambda_{c}$, determined via finite size scaling as discussed below, are also shown. The collapse of the $\overline{I}_{s_{0}}$ curves for different spatial sizes on top of each other starts precisely at $\lambda_{c}$.
Figure 4: Mobility edge obtained from $I_{s_{0}}$ via a finite size scaling study, in lattice units (top panel) and physical units (bottom panel). Red circles correspond to fits including data from all system sizes, and blue squares to fits exluding data from the smallest size, $L=40$. Different data points at the same $\beta$ correspond to different fitting intervals around the mobility edge.
@@
We fitted our data for $\overline{I}_{s_{0}}$ (with $\Delta\lambda=0.0075$) to a polynomial truncation of the scaling function Eq. (27), obviously setting $L=N_{s}$, and using the constrained fitting technique of Ref. [54]. This amounts in practice to minimizing an augmented $\chi^{2}_{\mathrm{aug}}=\chi^{2}+\chi^{2}_{\mathrm{prior}}$, where
@@
Our results for $\lambda_{c}$ are shown in Fig. 4, top panel. The first surprising aspect is that $\lambda_{c}$ exceeds the lowest Matsubara frequency $\lambda_{(0)}$. Localized modes then eat into the bulk of the spectrum, with disorder in the link variables being able to localize modes that one would expect to be extended, instead of modes localized on favorable gauge-field fluctuations becoming delocalized due to mixing with delocalized modes. The second, even more surprising aspect is that $\lambda_{c}$ in lattice units decreases with $\beta$, while in physical units it is compatible with a constant within 1%, see Fig. 4, bottom panel. This suggests that the mobility edge jumps from zero to a finite value at the critical value of the coupling, $\beta_{c}$, unexpectedly since the thermal transition is continuous.
@@
To check our surprising results for the $\beta$ dependence of $\lambda_{c}$, we estimated the mobility edge using a different approach, less rigorous but also less prone to fit instabilities due to the peculiar nature of the Anderson transition in this system. The idea is that the value of $\lambda$ at which $\overline{I}_{s_{0}}$ takes any prescribed value between the Poisson expectation and the critical value found at $\lambda_{c}$ will eventually tend to the mobility edge as the volume is increased. We then did a spline interpolation $I_{s_{0}}^{\mathrm{spline}}(\lambda)$ of the numerical data for each $\beta$ and $L$, and estimated $\lambda_{c}^{\mathrm{spline}}$ as the solution to $I_{s_{0}}^{\mathrm{spline}}(\lambda_{c}^{\mathrm{spline}}(\beta,L))=0.185$. The specific value on the right hand side of the equation was chosen safely above the (possibly $\beta$-dependent) critical value at the mobility edge, estimated by observing where the curves for the various volumes fall on top of each other at each $\beta$. We then obtained our alternative estimate of the mobility edge by extrapolating to infinite volume, $\lambda_{c}^{\mathrm{spline}}(\beta)=\lim_{L\to\infty}\lambda_{c}^{\mathrm{spline}}(\beta,L)$, assuming a linear dependence on $1/L$. We estimated the systematics from finite size effects by repeating the extrapolation after excluding the smallest volume. For $\beta=0.74$ the result for the largest volume is clearly off of a linear trend, and its inclusion in the fit leads to a very large $\chi^{2}/\mathrm{d.o.f.}$. We have then exluded it from the fit in the determination of the central value, while including the change due to its inclusion in the estimate of the error. In Fig. 6 the results are compared with the average $\lambda_{c}^{\mathrm{FSS}}$ of the determinations of $\lambda_{c}$ via finite size scaling shown in Fig. 4, with total error estimated adding in quadrature the average of the statistical error and the standard deviation over the set of determinations. The two estimates $\lambda_{c}^{\mathrm{FSS}}$ and $\lambda_{c}^{\mathrm{spline}}$ are in agreement within errors, although the alternative estimates is always smaller than the finite size scaling one. This is not entirely surprising, as $\lambda_{c}^{\mathrm{spline}}(\beta,L)$ always underestimates $\lambda_{c}$ by construction, and the linear extrapolation to infinite volume is probably neglecting sizeable effects. In any case, the important aspect is that the two estimates agree on the qualitative behavior of $\lambda_{c}$ as a function of $\beta$.
@@
The decrease of $\lambda_{c}$ with $\beta$ can be understood noticing that the width of the spectral pseudogap, where the density of modes is low, essentially coincides with the lowest Matsubara frequency, which in lattice units depends only on $N_{t}$ and so is fixed here. This makes the low end an effective edge of the spectrum, where disorder localizes the outermost modes first. Increasing $\beta$ at fixed $N_{t}$, the disorder decreases as the system gets more ordered, pushing the mobility edge towards the end of the spectrum, which in this case means reducing $\lambda_{c}$. The decrease in disorder is apparently compensated by the decrease of the lattice spacing $1/\beta=e^{2}a$, so that in physical units the mobility edge stays constant. We return on this point below in Sec. III.4. The difference with other theories studied so far is evident, and presumably due to the discrete nature of the gauge group. In the other cases where the dependence of $\lambda_{c}$ on $\beta$ was studied so far [6, 13, 14, 15, 16, 17, 18, 21], the gauge group was continuous, and in that case probably the spectral pseudogap does not depend only on $N_{t}$ but on $\beta$ as well.
@@ III.2 Center cluster percolation
To study the percolation properties of center clusters across the transition we performed a second set of simulations on lattices with $N_{s}=24,32,40,48,56,64,72,80$, accumulating between 60k and 135k configurations for the various volumes. For each configuration, after counting the number of active bonds, we identified all the center clusters, finding then the largest cluster and measuring its size in terms of the dual sites it touched. We then determined the concentration of active bonds $\alpha$, Eq. (13), and largest cluster density $\mathcal{S}$, Eq. (12), as functions of the lattice coupling $\beta$. Results for $\alpha(\beta)$ for various lattice sizes are shown in Fig. 7, and results for $\mathcal{S}(\beta)$ are shown in Fig. 8. The function $\alpha(\beta)$ is monotonic, and so invertible. At low $\beta$, corresponding to larger active-bond concentrations, $\mathcal{S}(\beta)$ depends mildly on the volume and tends to a finite value as the volume increases, signalling the presence of an infinite cluster.
@@
We then studied quantitively the volume scaling of the largest-cluster density $\mathcal{S}$ as a function of the active-bond concentration $\alpha$. To show that the system does indeed undergo a percolation transition, we performed a finite-size scaling analysis of $\mathcal{S}$ as a function of $\alpha$. We fitted our numerical data with the functional form
@@
The critical exponents clearly deviate from those of two-dimensional percolation (see Ref. [56]). This is probably due to the topological constraint on clusters mentioned above in Sec. II.3, that changes the universality class of the transition.
@@
Figure 10: Infinite-volume extrapolation of the critical coupling $\beta_{c}^{(p)}$ for the percolation transition, for $N_{t}=4$. Straight dashed lines correspond to linear extrapolations in $1/N_{s}^{2}$, each ranging only over the volumes used in the extrapolation. The filled black square is our final result for the extrapolation, including statistical and finite-size systematic errors, while the empty black square is the critical lattice coupling for deconfinement obtained in Ref. [34].
@@
Figure 12: Average rank in magnitude of the cluster $C_{\mathrm{max}}$ on which modes have maximal weight at $\beta=0.69$ (top panel), $\beta=0.72$ (center panel), in the confined phase, and $\beta=0.74$ (bottom panel), in the deconfined phase, for $N_{t}=4$ and various $N_{s}$. Dashed vertical lines mark the ends of the bulk, $\lambda_{(0)}$ and $\lambda_{(1)}$; the solid vertical line in the bottom panel marks the mobility edge $\lambda_{c}$ between low and bulk modes.
@@ III.3 Correlation between eigenmodes and clusters
In Fig. 12 we show the average rank in magnitude of the cluster $C_{\mathrm{max}}$ on which modes have maximal weight (see Sec. II.3), studied locally in spectrum. At $\beta=0.69$ (top panel), well within the confined phase, this is close to 1 for low and bulk modes, reaching slightly higher values for high modes. This in agreement with the modes’ localization properties: the delocalized low and bulk modes favor the largest cluster, which is the infinite cluster extended over the whole lattice discussed in the previous Subsection, while the localized high modes tend to favor smaller clusters, of lower rank, and more so as the system size increases. At $\beta=0.72$ (center panel), closer to the deconfined phase, low modes start showing a tendency to favor smaller clusters. At $\beta=0.74$ (bottom panel), well within the deconfined phase, the infinite cluster has disappeared, and the relation between the eigenvalue and the average rank of the corresponding $C_{\mathrm{max}}$ is far from straightforward, with the lowest modes not favoring the largest cluster, as one would naively expect, and marked peaks in the low-lying part of the spectrum (already visible, to a lower extent, already in the confined phase).
The cluster IPR, $\mathrm{IPR}_{\mathrm{clust}}$, Eq. (18), averaged locally in the spectrum according to Eq. (9), is shown for $\beta=0.69$ in the confined phase and $\beta=0.74$ in the deconfined phase in the top and center panels of Fig. 13, respectively. At $\beta=0.69$, both low and bulk modes spread on average on $1.5\div 2$ clusters, including the infinite cluster, as shown above, with a mild increase in the spreading as $N_{s}$ increases. Localized high modes instead typically lie on fewer clusters ($1\div 1.5$), without any apparent volume dependence of $\mathrm{IPR}_{\mathrm{clust}}$. This situation changes in the deconfined phase, where the infinite cluster is absent. For bulk modes, $\mathrm{IPR}_{\mathrm{clust}}\simeq 2\langle n_{\mathrm{clust}}\rangle^{-1}$ (see Fig. 13, bottom panel), i.e., they spread on about half of the $n_{\mathrm{clust}}$ (finite) clusters in a configuration; this also determines the marked volume dependence of $\mathrm{IPR}_{\mathrm{clust}}$, since $\langle n_{\mathrm{clust}}\rangle\propto V$, see Fig. 11. Low and high modes, instead, spread on about $1.5\div 3$ and $1\div 3$ clusters, respectively, with a mild volume dependence. This is again in agreement with the modes’ localization properties.
The maximal weight on a cluster $W_{\mathrm{max}}$, Eq. (16), again averaged locally in the spectrum, is shown for the same values of $\beta$ in the top and center panels of Fig. 14. In the confined phase this quantity does not show a big difference between low and bulk modes, and a mild volume dependence for both, reflecting the fact that these modes are delocalized on the infinite center cluster. High modes show a larger value and very little volume dependence, reflecting their localized nature. In the deconfined phase the difference between low and bulk modes becomes marked, and only bulk modes show a noticeable, $1/V$ volume dependence due to their delocalizing over a finite fraction of clusters (see Fig. 14, bottom panel).
@@
Figure 17: The quantity $\mathcal{A}_{\psi}$, measuring the correlation between islands and eigenmodes, for several values of $\beta$ both in the confined and deconfined phase, on a $28^{2}\times 4$ lattice. Vertical dashed lines correspond to the bulk ends $\lambda_{(0)}$ and $\lambda_{(1)}$.

### v2

@@ II.5 Sea/islands picture and Dirac-Anderson Hamiltonian
For the staggered Dirac operator the sea/islands picture briefly sketched in the Introduction can be understood more precisely in the language of the Dirac-Anderson Hamiltonian [31, 32, 24, 20]. In ($d+1$) dimensions, in the basis where the temporal part of the staggered operator is diagonal, the “Hamiltonian” $-iD^{\mathrm{stag}}$ reads $H^{\rm DA}=-i\Omega^{\dagger}D^{\mathrm{stag}}\Omega$, where $\Omega$ is a suitable unitary change of basis and [20]
@@
${\cal A}_{\psi}(\lambda)\equiv\frac{\left\langle\sum_{l}\delta(\lambda-\lambda_{l})\sum_{\vec{x}}{\cal A}(\vec{x})\sum_{t=0}^{N_{t}-1}|\psi_{l}(t,\vec{x})|^{2}\right\rangle}{\left\langle\sum_{l}\delta(\lambda-\lambda_{l})\right\rangle\langle{\cal A}(\vec{x})\rangle}\,.$ (30)
@@ III.1 Mobility edge
In Figs. 1 and 2 we show the normalized spectral density $\rho(\lambda)$, Eq. (10), and the average mode size $\overline{\mathrm{IPR}^{-1}}$, see under Eq. (20), along the low-lying spectrum. The spectral density is small but nonzero near $\lambda=0$, and starts to increase rapidly in the bulk, above $\lambda_{(0)}$. For low modes the mode size is essentially volume independent, indicating localization. Peak structures are present in $\rho$ at specific values of $\lambda$ below $\lambda_{(0)}$, where one correspondingly finds dips in the mode size. We will return on this feature in Sec. III.5.
We then studied the statistical properties of the unfolded spectrum. Since degenerate eigenvalues are absent in the spectral region and for the volumes under consideration, we have obtained the unfolded eigenvalues numerically by sorting in ascending order all the eigenvalues on all the configurations for a given choice of $\beta$ and $N_{s}$, and assigning to each eigenvalue its rank divided by the number of configurations. We then used them to numerically estimate $\overline{I}_{s_{0}}(\lambda,N_{s})$, Eq. (25), as explained above, and studied its volume dependence. The behavior expected for a BKT-type Anderson transitions is clearly visible in Fig. 3, where we show $\overline{I}_{s_{0}}$ for different spatial volumes at $\beta=0.734$ and $\beta=0.7425$ (here $\Delta\lambda=0.0125$). While the low modes tend towards Poissonian statistics as the system size increases, higher up in the spectrum the spectral statistics does not change with $N_{s}$. The lowest lattice Matsubara frequency, $\lambda_{(0)}$, and the mobility edge, $\lambda_{c}$, determined via finite size scaling as discussed later, are also shown. The collapse of the $\overline{I}_{s_{0}}$ curves for different spatial sizes on top of each other starts precisely at $\lambda_{c}$.
Figure 4: Mobility edge obtained from $I_{s_{0}}$ via a finite size scaling study, in lattice units (top panel) and physical units (bottom panel). Red circles correspond to fits including data from all system sizes, and blue squares to fits excluding data from the smallest size, $L=40$. Different data points at the same $\beta$ correspond to different fitting intervals around the mobility edge.
@@
We fitted our data for $\overline{I}_{s_{0}}$ (with $\Delta\lambda=0.0075$) to a polynomial truncation of the scaling function Eq. (27), obviously setting $L=N_{s}$, and using the constrained fitting technique of Ref. [77]. This amounts in practice to minimizing an augmented $\chi^{2}$, i.e., $\chi^{2}_{\mathrm{aug}}=\chi^{2}+\chi^{2}_{\mathrm{prior}}$, where
@@
Our results for $\lambda_{c}$ are shown in Fig. 4, top panel. The first surprising aspect is that $\lambda_{c}$ exceeds the lowest lattice Matsubara frequency $\lambda_{(0)}$. Localized modes then eat into the bulk of the spectrum, with disorder in the link variables being able to localize modes that one would expect to be extended, instead of modes localized on favorable gauge-field fluctuations becoming delocalized due to mixing with delocalized modes. The second, even more surprising aspect is that $\lambda_{c}$ in lattice units decreases with $\beta$, while in physical units it is compatible with a constant within 1%, see Fig. 4, bottom panel. This suggests that the mobility edge jumps from zero to a finite value at the critical value of the coupling, $\beta_{c}$, unexpectedly since the thermal transition is continuous.
@@
To check our surprising results for the $\beta$ dependence of $\lambda_{c}$, we estimated the mobility edge using a different approach, less rigorous but also less prone to fit instabilities due to the peculiar nature of the Anderson transition in this system. The idea is that the value of $\lambda$ at which $\overline{I}_{s_{0}}$ takes any prescribed value between the Poisson expectation and the critical value found at $\lambda_{c}$ will eventually tend to the mobility edge as the volume is increased. We then did a spline interpolation $I_{s_{0}}^{\mathrm{spline}}(\lambda)$ of the numerical data for each $\beta$ and $L$, and estimated $\lambda_{c}^{\mathrm{spline}}$ as the solution to $I_{s_{0}}^{\mathrm{spline}}(\lambda_{c}^{\mathrm{spline}}(\beta,L))=0.185$. The specific value on the right-hand side of the equation was chosen safely above the (possibly $\beta$ dependent) critical value at the mobility edge, estimated by observing where the curves for the various volumes fall on top of each other at each $\beta$. We then obtained our alternative estimate of the mobility edge by extrapolating to infinite volume, $\lambda_{c}^{\mathrm{spline}}(\beta)=\lim_{L\to\infty}\lambda_{c}^{\mathrm{spline}}(\beta,L)$, assuming a linear dependence on $1/L$. We estimated the systematics from finite size effects by repeating the extrapolation after excluding the smallest volume. For $\beta=0.74$ the result for the largest volume is clearly off of a linear trend, and its inclusion in the fit leads to a very large $\chi^{2}/\mathrm{d.o.f.}$. We have then exluded it from the fit in the determination of the central value, while including the change due to its inclusion in the estimate of the error. In Fig. 6 the results are compared with the average $\lambda_{c}^{\mathrm{FSS}}$ of the determinations of $\lambda_{c}$ via finite size scaling shown in Fig. 4, with total error estimated adding in quadrature the average of the statistical error and the standard deviation over the set of determinations. The two estimates $\lambda_{c}^{\mathrm{FSS}}$ and $\lambda_{c}^{\mathrm{spline}}$ are in agreement within errors, although the alternative estimate is always smaller than the finite size scaling one. This is not entirely surprising, as $\lambda_{c}^{\mathrm{spline}}(\beta,L)$ always underestimates $\lambda_{c}$ by construction, and the linear extrapolation to infinite volume is probably neglecting sizable effects. In any case, the important aspect is that the two estimates agree on the qualitative behavior of $\lambda_{c}$ as a function of $\beta$.
@@
The decrease of $\lambda_{c}$ with $\beta$ can be understood noticing that the width of the spectral pseudogap, where the density of modes is low, essentially coincides with the lowest lattice Matsubara frequency, $\lambda_{(0)}$, which depends only on $N_{t}$ and so is fixed here. This makes the low end an effective edge of the spectrum, where disorder localizes the outermost modes first. Increasing $\beta$ at fixed $N_{t}$, the disorder decreases as the system gets more ordered, pushing the mobility edge towards the end of the spectrum, which in this case means reducing $\lambda_{c}$. The decrease in disorder is apparently compensated by the decrease of the lattice spacing $1/\beta=e^{2}a$, so that in physical units the mobility edge stays constant. We return on this point below in Sec. III.4. The difference with other theories studied so far is evident, and presumably due to the discrete nature of the gauge group. In the other cases where the dependence of $\lambda_{c}$ on $\beta$ was studied so far [6, 13, 14, 15, 16, 17, 18, 21, 22], the gauge group was continuous, and in that case probably the spectral pseudogap does not depend only on $N_{t}$ but on $\beta$ as well.
@@ III.2 Center cluster percolation
To study the percolation properties of center clusters across the transition we performed a second set of simulations on lattices with $N_{s}=24,32,40,48,56,64,72,80$, accumulating between 60k and 135k configurations for the various volumes. For each configuration, after counting the number of active bonds, we identified all the center clusters, finding then the largest cluster and measuring its size in terms of the dual sites it touched. We then determined the concentration of active bonds $\alpha$, Eq. (13), and largest cluster density $\mathcal{S}$, Eq. (12), as functions of the lattice coupling $\beta$. Results for $\alpha(\beta)$ for various lattice sizes are shown in Fig. 7, and results for $\mathcal{S}(\beta)$ are shown in Fig. 8. The function $\alpha(\beta)$ is monotonic, and so invertible. At low $\beta$, corresponding to larger active-bond concentrations, $\mathcal{S}(\beta)$ depends mildly on the volume and tends to a finite value as the volume increases, signaling the presence of an infinite cluster.
@@
We then studied quantitatively the volume scaling of the largest-cluster density $\mathcal{S}$ as a function of the active-bond concentration $\alpha$. To show that the system does indeed undergo a percolation transition, we performed a finite-size scaling analysis of $\mathcal{S}$ as a function of $\alpha$. We fitted our numerical data with the functional form
@@
The critical exponents clearly deviate from those of two-dimensional bond percolation, $\nu=4/3$ and $\beta/\nu=5/48$ (see Refs. [60, 61, 62, 63]), in spite of their expected universality (see the proof in Ref. [63] for square, triangular, and hexagonal lattices). This is probably due to the topological constraint imposed on clusters, mentioned above in Sec. II.3, and possibly to the presence of an additional dimension (albeit of finite extent), that change the universality class of the transition.
@@
Figure 10: Infinite-volume extrapolation of the critical coupling $\beta_{c}^{(p)}$ for the percolation transition, for $N_{t}=4$. The three dashed straight lines correspond to linear extrapolations in $1/N_{s}^{2}$, each ranging only over the volumes used in the extrapolation. The filled black square is our final result for the extrapolation, including statistical and finite-size systematic errors, while the empty black square is the critical lattice coupling for deconfinement obtained in Ref. [34].
@@
Figure 12: Average rank in magnitude of the cluster $C_{\mathrm{max}}$ on which modes have maximal weight at $\beta=0.69$ (top panel) and $\beta=0.72$ (center panel), in the confined phase, and $\beta=0.74$ (bottom panel), in the deconfined phase, for $N_{t}=4$ and various $N_{s}$. Dashed vertical lines mark the ends of the bulk, $\lambda_{(0)}$ and $\lambda_{(1)}$; the solid vertical line in the bottom panel marks the mobility edge $\lambda_{c}$ between low and bulk modes.
@@ III.3 Correlation between eigenmodes and clusters
In Fig. 12 we show the average rank in magnitude of the cluster $C_{\mathrm{max}}$ on which modes have maximal weight (see Sec. II.3), studied locally in spectrum. At $\beta=0.69$ (top panel), well within the confined phase, this is close to 1 for low and bulk modes, reaching slightly higher values for high modes. This is in agreement with the modes’ localization properties: the delocalized low and bulk modes favor the largest cluster, which is the infinite cluster extended over the whole lattice discussed in the previous subsection, while the localized high modes tend to favor smaller clusters, of lower rank, and more so as the system size increases. At $\beta=0.72$ (center panel), closer to the deconfined phase, low modes start showing a tendency to favor smaller clusters. At $\beta=0.74$ (bottom panel), well within the deconfined phase, the infinite cluster has disappeared, and the relation between the eigenvalue and the average rank of the corresponding $C_{\mathrm{max}}$ is far from straightforward, with the lowest modes not favoring the largest cluster, as one would naively expect, and marked peaks in the low-lying part of the spectrum (already visible, to a lower extent, already in the confined phase).
[…]
The maximal weight on a cluster $W_{\mathrm{max}}$, Eq. (16), again averaged locally in the spectrum, is shown for the same values of $\beta$ in the top and center panels of Fig. 14. In the confined phase this quantity does not show a big difference between low and bulk modes, and a mild volume dependence for both, reflecting the fact that these modes are delocalized on the infinite center cluster. High modes show a larger value and very little volume dependence, reflecting their localized nature. In the deconfined phase the difference between low and bulk modes becomes marked, and only bulk modes show a noticeable, $1/V$ volume dependence due to their delocalizing over a finite fraction of clusters (see Fig. 14, bottom panel). For the lowest and highest modes, most of the mode weight is found on a single finite cluster even as the volume increases; closer to the mobility edges [^3] We did not determine precisely the position of the mobility edge separating delocalized bulk and localized high modes, but all observables point at it being found near the high end of the bulk. the maximal fraction of weight on a cluster goes below $0.5$, but it still seems to converge to a finite value in the large-volume limit.
@@
The results of this subsection show that the localized low modes appearing in the deconfined phase (and the localized high modes present both in the confined and in the deconfined phase) have a clear tendency to localize on one, or at most a few center clusters. This shows that the gauge field structures related to the confining properties of the theory coincide with the localization centers for these modes.
@@
Figure 17: The quantity $\mathcal{A}_{\psi}$, measuring the correlation between islands and eigenmodes, for several values of $\beta$ both in the confined and deconfined phase, on a $4\times 28^{2}$ lattice. Vertical dashed lines correspond to the bulk ends $\lambda_{(0)}$ and $\lambda_{(1)}$.